Suppose that 75% of all dialysis patients will survive for at least 5 years.

In a simple random sample of 100 new dialysis patients, what is the probability that the proportion surviving for at least five years will exceed 80%, rounded to 5 decimal places?
Answer #1
for normal distribution z score =(p̂-p)/σp
here population proportion=     p= 0.750
sample size       =n= 100
std error of proportion=σp=√(p*(1-p)/n)= 0.0433
probability = P(X>0.8) = P(Z>1.15)= 1-P(Z<1.15)= 1-0.87589= 0.12411